Write down 𝑇3(π‘₯), 𝑇4(π‘₯), π‘Žπ‘›π‘‘ 𝑇5(π‘₯) for the Taylor series of 𝑓(π‘₯) = ln (3 + 4π‘₯) about π‘₯ = 0


1
Expert's answer
2022-03-03T13:50:37-0500
"f(x)=\ln(3+4x)"


"f(0)=\ln(3+4(0))=\ln3"

"f'(x)=\dfrac{4}{3+4x}"

"f'(0)=\dfrac{4}{3+4(0)}=\dfrac{4}{3}"


"f''(x)=-\dfrac{16}{(3+4x)^2}"

"f''(0)=-\dfrac{16}{(3+4(0)^2}=-\dfrac{16}{9}"

"f'''(x)=\dfrac{128}{(3+4x)^3}"

"f'''(0)=\dfrac{128}{(3+4(0))^3}=\dfrac{128}{27}"

"f^{(4)}(x)=-\dfrac{1536}{(3+4x)^4}"

"f^{(4)}(0)=-\dfrac{1536}{(3+4(0))^4}=-\dfrac{512}{27}"

"f^{(5)}(x)=\dfrac{24576}{(3+4x)^5}"

"f^{(5)}(0)=\dfrac{24576}{(3+4(0))^5}=\dfrac{8192}{81}"

"T_3(x)=\ln3+\dfrac{4}{3}x-\dfrac{8}{9}x^2+\dfrac{64}{81}x^3"



"T_4(x)=\ln3+\dfrac{4}{3}x-\dfrac{8}{9}x^2+\dfrac{64}{81}x^3-\dfrac{64}{81}x^4"



"T_5(x)=\ln3+\dfrac{4}{3}x-\dfrac{8}{9}x^2+\dfrac{64}{81}x^3-\dfrac{64}{81}x^4"

"+\dfrac{1024}{1215}x^5"